the length of a rectanglar garden is 5m greater than the width. the area is 66^2. find the dim. of the garden.
Answer by checkley77(12844)

L=W+5
LW=66
W(W+5)=66
W^2+5W-66=0
(W+11)(W-6)=0
W-6=0
W=6 M. IS THE WIDTH.
L=6+5=11 M. IS THE LENGTH.
PROOF:
6*11=66
66=66
---------------------------------------
20b^2-64+64
20B^2 ANS.